how do I Calculate the rate difference and the rate ratio from the following data  

# of Patients

Male

Female

Critical Care

6

2

Healthy

48

39
Answer #1

Rate difference of male and female

For critical care

Rate difference = { (6-2) / Average of 6 and 2} * 100 =

=|6−2| / [(6+2/2]×100

=4/4*100

=1×100=1×100

=100%difference

For Healthy

Rate difference = { (48-39) / Average of 48 and 39 } *100

=(9 / 43.5 ) *100

= 20.69% difference

Rate Ratio

of Patients

Male

Female

Critical Care

a = 6

b = 2

Healthy

c=48

d= 39

Rate ratio

For Male

Rate ratio = 6/48 = 1/8

For Female

Rate ratio = 2/39